# T-Retrievability: A Topic-Focused Approach to Measure Fair Document Exposure in Information Retrieval

### v1

@@ Abstract.
Retrievability of a document is a collection-based statistic that measures its expected (reciprocal) rank of being retrieved within a specific rank cut-off. A collection with uniformly distributed retrievability scores across documents is an indicator of fair document exposure. While retrievability scores have been used to quantify the fairness of exposure for a collection, in our work, we use the distribution of retrievability scores to measure the exposure bias of retrieval models. We hypothesise that an uneven distribution of retrievability scores across the entire collection may not accurately reflect exposure bias but rather indicate variations in topical relevance. As a solution, we propose a topic-focused localised retrievability measure, which we call T-Retrievability (topic-retrievability), which first computes retrievability scores over multiple groups of topically-related documents, and then aggregates these localised values to obtain the collection-level statistics. Our analysis using this proposed T-Retrievability measure uncovers new insights into the exposure characteristics of various neural ranking models. The findings suggest that this localised measure provides a more nuanced understanding of exposure fairness, offering a more reliable approach for assessing document accessibility in IR systems.
@@ 1. Introduction
Supervised information retrieval (IR) models based on pre-trained transformer architectures, such as ColBERT (Khattab and Zaharia, 2020) and Mono-T5 (Nogueira et al., 2020), have demonstrated superior relevance ranking performance compared to their unsupervised counterparts, such as BM25 (Robertson et al., 1994). However, this paradigm shift necessitates the consideration of additional dimensions of effectiveness, including those of efficiency (Bruch et al., 2024; Chang et al., 2024), explainability (Verma and Ganguly, 2019; Pandian et al., 2024; Sen et al., 2020), and exposure fairness (Singh and Joachims, 2018; Heuss et al., 2022; Patro et al., 2022). In this study, we specifically examine the exposure fairness criterion of search systems, which has significant societal implications. For instance, ensuring a balanced distribution of topically relevant documents from diverse sources and ideological viewpoints is essential for preserving the integrity of information ecosystems (Patro et al., 2022), fostering an informed and unbiased public opinion, and supporting democratic decision-making (Kulshrestha et al., 2017; Krause et al., 2024; Abolghasemi et al., 2024).
@@
The data-driven learning of relevance in neural models makes it challenging to identify and mitigate latent biases present in the training data, which can manifest as algorithmic preferences leading to exposure biases (Patro et al., 2022; Liu et al., 2023; Musto et al., 2021). These exposure biases are expected ranks of retrieved documents computed over a large set of queries that several users may execute on a collection. In the case of queries with similar information needs and overlapping sets of relevant documents, an IR model that aims to ensure exposure fairness should distribute the rankings of these documents more evenly by shuffling their positions across queries (Diaz et al., 2020; Ekstrand et al., 2023). Assessing whether IR models achieve this fairness objective is crucial for understanding and addressing potential biases in retrieval systems. We restrict the scope of our investigation of exposure bias to only document collections without categorical meta-data attributes, e.g., gender, polarity etc., as studied in (Rekabsaz et al., 2021).
@@
Previous research has introduced the concept of retrievability as a measure of document accessibility (Azzopardi and Vinay, 2008; Wilkie and Azzopardi, 2014, 2015, 2016). This collection-based statistic quantifies a document’s expected (reciprocal) rank of retrieval within a predefined rank cut-off, averaged over a sufficiently large and topically diverse set of queries. While retrievability analysis has traditionally been used to assess document accessibility across different collections (Wilkie and Azzopardi, 2014), more recent studies have applied it to examine exposure bias across various IR models (Abolghasemi et al., 2023). However, two key considerations must be taken into account when using retrievability for such analyses. First, most existing studies, including those evaluating exposure fairness in IR models (Abolghasemi et al., 2023), rely on synthetic queries generated via word $n$-gram sampling. Synthetic queries used in prior research predominantly consist of unigrams or bigrams, whereas real-world user queries tend to be longer on average. Second, non-uniform retrievability values often arise due to variations in relevance priors, where certain high-quality documents are inherently more likely to be relevant than others. In such cases, disparities in retrievability should be accounted for by aggregating locally computed retrievability measures over groups of similar information needs.
@@ 2. Proposed Measure for Exposure Fairness
As per the originally proposed definition of retrievability (Azzopardi and Vinay, 2008), its value for a document $D\in\pazocal{C}$ (a collection of documents) indicates the likelihood of the document to be retrieved within a specified cut-off rank. As such, the retrievability value of a document depends on the retrieval model used to induce the ranking itself, and a cut-off rank is usually set to a small number, such as 10, indicating the size of a typical search results page (Kelly and Azzopardi, 2015). More formally, $r(D,\pazocal{C},\pazocal{Q},\theta,k)=\frac{1}{|\pazocal{Q}|}\sum_{Q\in\pazocal{Q}}\mathbb{I}[\rho(D;Q,\theta)\leq k]$, where $D$ is a document of the collection $\pazocal{C}$, $\pazocal{Q}$ denotes a set of queries, $\theta:D\times Q\mapsto\mathbb{R}$ denotes an IR model that assigns scores to documents thereby inducing a rank for $D$ denoted as $\rho(D;Q,\theta)$, which is a function of both a particular query $Q$ and the IR model $\theta$. The parameter $k$ denotes a cut-off rank and $\mathbb{I}[X]$ denotes an indicator function which is 1 if the condition $X$ is true, otherwise 0.
@@
In this paper, we work with a slightly modified version of the retrievability measure, where we remove the dependence of the cut-off rank $k$. More specifically, instead of computing the likelihood of a document to be retrieved within a top-$k$, we rather compute the expected rank of a document (Sinha et al., 2023). In particular, to interpret retrievability as a score, i.e., the higher the better, we work with reciprocals of ranks instead of the ranks themselves. Formally,
@@
The bias in information access is computed by the Gini coefficient (Gini, 1921) over the retrievability distribution of each document in the collection - a high value of the coefficient indicating a high disparity between document accessibility.
@@
An adequately large set of representative queries $\pazocal{Q}$ is used to measure the accessibility of a document in a collection $\pazocal{C}$ (Equation 1). All previous work on retrievability analysis (Azzopardi and Vinay, 2008; Wilkie and Azzopardi, 2014; Azzopardi, 2015; Wilkie and Azzopardi, 2016; Ganguly et al., 2016) uses simulation to create this set $\pazocal{Q}$ - specifically via sampling words and their bi-grams from the collection. The $r(D)$ values can thus be affected by sampling biases resulting from the simulation. Instead, in our work, we conduct our experiments on real queries formulated by humans. Specifically, we employ the MS MARCO dev set as $\pazocal{Q}$ comprising a total of over 100K queries (collected from the Bing search log (Bajaj et al., 2016)).
@@
A likely reason why real-world query logs are often avoided in retrievability studies (Azzopardi and Vinay, 2008; Wilkie and Azzopardi, 2014; Azzopardi, 2015; Wilkie and Azzopardi, 2016; Ganguly et al., 2016) is that they tend to be biased toward specific topics or user interests. However, employing a sufficiently large number of such human-formulated queries from search logs, as in our work, can mitigate this bias, as the topical skew diminishes when aggregated across a broader query set.
@@
Previous research has shown that the prior distribution of relevance in a collection is usually non-uniform (Chang et al., 2024), meaning that some documents due to their inherent characteristics, such as their lengths (Singhal et al., 2017), citation counts (Peng et al., 2008) etc., are more likely to be relevant for a higher number of queries than others. However, this means that this non-uniformity in relevance-priors is also likely to be manifested as non-uniformities in the collection-level retrievability scores (see Equation 1) computed for a retrieval model that effectively models relevance, thereby falsely penalising the model.
@@
To address the influence of relevance priors on a collection-level document accessibility measure, such as retrievability, we propose to measure document accessibility for groups of topically related queries. This idea aligns with the per-query analysis of retrievability over a pool of retrieved documents as studied in (Wilkie and Azzopardi, 2016). This idea also aligns with the existing research on measuring exposure fairness for groups defined by document metadata (attribute value) combinations, such as demographics (Rekabsaz et al., 2021), location (Ekstrand et al., 2023), and polarities (Kulshrestha et al., 2017). In our setup, instead of measuring disparities across groups of attribute values, we investigate disparities in document accessibility.
@@
For a fine-grained computation of retrievability scores, we partition the query set $\pazocal{Q}$ (in our experiments, the MS MARCO development queries) into $K$ clusters (Lupart et al., 2023). Specifically, we apply K-means clustering using two types of document vector representations: sparse (lexical) embeddings and dense (semantic) embeddings. These representations yield query groupings with distinct characteristics—one emphasizing precision and the other recall. The lexical approach is more conservative, grouping queries into the same cluster only when they share exact term matches, thereby achieving higher precision. In contrast, the semantic approach expands clusters based on term-level semantics, which increases recall but introduces a greater risk of reduced precision within each cluster.
For the sparse representation, we employ the standard TF–IDF vectorization of each query. For the dense representation, we use the [CLS] token’s vector computed with the Sentence-BERT (SBERT) model (Reimers and Gurevych, 2019). To ensure fair comparison across neural models, we adopt the all-MiniLM-L6-v2 variant of SBERT – a model pre-trained to capture semantic similarities (Bowman et al., 2015).
@@
(3) $\operatorname*{G}[(D,\pazocal{C},\pazocal{Q},\theta),\oplus]=\oplus_{i=1}^{K}\operatorname*{G}[r(D,\pazocal{C},\pazocal{Q}_{i},\theta)],$
where $\bar{G}[\cdot]$ represents the Gini coefficient of a set of localised retrievability values, $\oplus\in\{\min,\operatorname*{avg},\max\}$ denotes an aggregation operator. We call this new collection-level measure of document accessibility Equation 3 by the name Topical-Retrievability, abbreviated as T-Retrievability [^1] The implementation is available at https://github.com/XuejunChang/T-Retrievability .
@@ 3. Experiment Setup
It is worth noting that our proposed localized retrievability scores depend on the granularity of the topical relatedness of queries, i.e., too coarse topics may lead to relevance prior biases, as is expected to be the case with the standard collection-level retrievability (Azzopardi and Vinay, 2008), whereas too fine-grained topics may fail to capture a collection-level accessibility (Wilkie and Azzopardi, 2016). Our third research question is thus:
@@
All our experiments corresponding to retrievability analysis and evaluating the relevance of different rankers are conducted based on the MS MARCO dev dataset, which is comprised of 101,093 queries and their associated relevant documents. The underlying collection of documents used is the MS MARCO passage collection, comprised of over 8.8 million passages (Bajaj et al., 2016).
@@
  • BM25 (Robertson et al., 1994): BM25 is a member of the sparse family of IR models; the term weighting scheme is a combination of the relative importance of terms within documents, term informativeness and document lengths.
  • SPLADE (Formal et al., 2021): SPLADE, which belongs to the learned-sparse family of rankers, combines the posterior likelihoods of the masked language model (MLM) objective of BERT, and the noise contrastive estimation likelihoods to derive the token weights.
  • TCT-ColBERT (abbreviated in Table 1 as ‘TCT’) (Lin et al., 2021): This is a dense end-to-end model distilled from ColBERT that learns effective document encoded vectors. An approximate nearest neighbour search on these embedded representations has been reported to yield better performance than the [CLS] pooling or mean pooling of the ColBERT model. In our experiments, we use the pyterrier_dr library to construct the dense index.
  • BM25¿¿TCT-ColBERT (abbreviated in Table 1 as ‘BM25¿¿TCT’) (Lin et al., 2021): the retrieve-and-rerank version of TCT-ColBERT, where we rerank the BM25 top-100 documents with the TCT-ColBERT scores for each query-document pair.
  • BM25¿¿Mono-T5 (abbreviated in Table 1 as ‘Mono-T5’) (Nogueira et al., 2020): A retrieve-and-rerank class of model, where the top 100 documents retrieved by BM25 are subsequently reranked using a cross-encoder model MonoT5, which assigns relevance scores to each query-document pair.
@@
As evaluation metrics for document exposure bias, we employ the Gini coefficient of the retrievability scores (lower scores indicating fair document exposure). As the measure for relevance, we employ the common measure of nDCG@10 (Järvelin and Kekäläinen, 2002) and MAP (Buckley and Voorhees, 2000).

### v2

@@ Abstract.
Retrievability of a document is a collection-based statistic that measures its expected (reciprocal) rank of being retrieved within a specific rank cut-off. A collection with uniformly distributed retrievability scores across documents is an indicator of fair document exposure. While retrievability scores have been used to quantify the fairness of exposure for a collection, in our work, we use the distribution of retrievability scores to measure the exposure bias of retrieval models. We hypothesise that an uneven distribution of retrievability scores across the entire collection may not accurately reflect exposure bias but rather indicate variations in topical relevance. As a solution, we propose a topic-focused localised retrievability measure, which we call T-Retrievability (topic-retrievability), which first computes retrievability scores over multiple groups of topically-related documents, and then aggregates these localised values to obtain the collection-level statistics. Our analysis using this proposed T-Retrievability measure uncovers new insights into the exposure characteristics of various neural ranking models. The findings suggest that this localised measure provides a more nuanced understanding of exposure fairness,
@@ 1. Introduction
Supervised information retrieval (IR) models based on pre-trained transformer architectures, such as ColBERT (colbert) and Mono-T5 (monot5), have demonstrated superior relevance ranking performance compared to their unsupervised counterparts, such as BM25 (BM25_RobertsonWJHG94). However, this paradigm shift necessitates the consideration of additional dimensions of effectiveness, including those of efficiency (BruchNRV24; chang2024neural), explainability (VermaG19; DBLP:conf/ecir/PandianGM24; DBLP:conf/sigir/SenGVJ20), and exposure fairness (singh2018fairness; FairnessExposure2022; patro2022fair). In this study, we specifically examine the exposure fairness criterion of search systems, which has significant societal implications. For instance, ensuring a balanced distribution of topically relevant documents from diverse sources and ideological viewpoints is essential for preserving the integrity of information ecosystems (patro2022fair), fostering an informed and unbiased public opinion, and supporting democratic decision-making (opinion; krause2024relevance; abolghasemi2024measuring).
@@
The data-driven learning of relevance in neural models makes it challenging to identify and mitigate latent biases present in the training data, which can manifest as algorithmic preferences leading to exposure biases (patro2022fair; Mitigating_Popularity_Bias2023; musto2021fairness). These exposure biases are expected ranks of retrieved documents computed over a large set of queries that several users may execute on a collection. In the case of queries with similar information needs and overlapping sets of relevant documents, an IR model that aims to ensure exposure fairness should distribute the rankings of these documents more evenly by shuffling their positions across queries (Fernando_2020; Ekstrand2023). Assessing whether IR models achieve this fairness objective is crucial for understanding and addressing potential biases in retrieval systems. We restrict the scope of our investigation of exposure bias to only document collections without categorical meta-data attributes, e.g., gender, polarity etc., as studied in (rekabsaz2021societal).
@@
Previous research has introduced the concept of retrievability as a measure of document accessibility (leifRetrievabilityAnevaluationMeasure; retr-analysis-cikm2014; wilkie2015retrievability; Topical_Retrievability2016). This collection-based statistic quantifies a document’s expected (reciprocal) rank of retrieval within a predefined rank cut-off, averaged over a sufficiently large and topically diverse set of queries. While retrievability analysis has traditionally been used to assess document accessibility across different collections (retr-analysis-cikm2014), more recent studies have applied it to examine exposure bias across various IR models (synth-queries). However, two key considerations must be taken into account when using retrievability for such analyses. First, most existing studies, including those evaluating exposure fairness in IR models (synth-queries), rely on synthetic queries generated via word $n$-gram sampling. Synthetic queries used in prior research predominantly consist of unigrams or bigrams, whereas real-world user queries tend to be longer on average. Second, non-uniform retrievability values often arise due to variations in relevance priors, where certain high-quality documents are inherently more likely to be relevant than others. In such cases, disparities in retrievability should be accounted for by aggregating locally computed retrievability measures over groups of similar information needs.
@@ 2. Proposed Measure for Exposure Fairness
As per the originally proposed definition of retrievability (leifRetrievabilityAnevaluationMeasure), its value for a document $D\in\pazocal{C}$ (a collection of documents) indicates the likelihood of the document to be retrieved within a specified cut-off rank. As such, the retrievability value of a document depends on the retrieval model used to induce the ranking itself, and a cut-off rank is usually set to a small number, such as 10, indicating the size of a typical search results page (serpsize). More formally, $r(D,\pazocal{C},\pazocal{Q},\theta,k)=\frac{1}{|\pazocal{Q}|}\sum_{Q\in\pazocal{Q}}\mathbb{I}[\rho(D;Q,\theta)\leq k]$, where $D$ is a document of the collection $\pazocal{C}$, $\pazocal{Q}$ denotes a set of queries, $\theta:D\times Q\mapsto\mathbb{R}$ denotes an IR model that assigns scores to documents thereby inducing a rank for $D$ denoted as $\rho(D;Q,\theta)$, which is a function of both a particular query $Q$ and the IR model $\theta$. The parameter $k$ denotes a cut-off rank and $\mathbb{I}[X]$ denotes an indicator function which is 1 if the condition $X$ is true, otherwise 0.
@@
In this paper, we work with a slightly modified version of the retrievability measure, where we remove the dependence of the cut-off rank $k$. More specifically, instead of computing the likelihood of a document to be retrieved within a top-$k$, we rather compute the expected rank of a document (sinha2023findability). In particular, to interpret retrievability as a score, i.e., the higher the better, we work with reciprocals of ranks instead of the ranks themselves. Formally,
@@
The bias in information access is computed by the Gini coefficient (gini1921measurement) over the retrievability distribution of each document in the collection - a high value of the coefficient indicating a high disparity between document accessibility.
@@
An adequately large set of representative queries $\pazocal{Q}$ is used to measure the accessibility of a document in a collection $\pazocal{C}$ (Equation 1). All previous work on retrievability analysis (leifRetrievabilityAnevaluationMeasure; retr-analysis-cikm2014; Retrievability_Azzo2015; Topical_Retrievability2016; retrievability_Ganguly2016) uses simulation to create this set $\pazocal{Q}$ - specifically via sampling words and their bi-grams from the collection. The $r(D)$ values can thus be affected by sampling biases resulting from the simulation. Instead, in our work, we conduct our experiments on real queries formulated by humans. Specifically, we employ the MS MARCO dev set as $\pazocal{Q}$ comprising a total of over 100K queries (collected from the Bing search log (bajaj2016ms)).
@@
A likely reason why real-world query logs are often avoided in retrievability studies (leifRetrievabilityAnevaluationMeasure; retr-analysis-cikm2014; Retrievability_Azzo2015; Topical_Retrievability2016; retrievability_Ganguly2016) is that they tend to be biased toward specific topics or user interests. However, employing a sufficiently large number of such human-formulated queries from search logs, as in our work, can mitigate this bias, as the topical skew diminishes when aggregated across a broader query set.
@@
Previous research has shown that the prior distribution of relevance in a collection is usually non-uniform (chang2024neural), meaning that some documents due to their inherent characteristics, such as their lengths (doclength-pivot), citation counts (docprior-pagerank) etc., are more likely to be relevant for a higher number of queries than others. However, this means that this non-uniformity in relevance-priors is also likely to be manifested as non-uniformities in the collection-level retrievability scores (see Equation 1) computed for a retrieval model that effectively models relevance, thereby falsely penalising the model.
@@
To address the influence of relevance priors on a collection-level document accessibility measure, such as retrievability, we propose to measure document accessibility for groups of topically related queries. This idea aligns with the per-query analysis of retrievability over a pool of retrieved documents as studied in (Topical_Retrievability2016). This idea also aligns with the existing research on measuring exposure fairness for groups defined by document metadata (attribute value) combinations, such as demographics (rekabsaz2021societal), location (Ekstrand2023), and polarities (opinion). In our setup, instead of measuring disparities across groups of attribute values, we investigate disparities in document accessibility.
@@
For a fine-grained computation of retrievability scores, we partition the query set $\pazocal{Q}$ (in our experiments, the MS MARCO development queries) into $K$ clusters (msmarco-clusters). Specifically, we apply K-means clustering using two types of document vector representations: sparse (lexical) embeddings and dense (semantic) embeddings. These representations yield query groupings with distinct characteristics—one emphasizing precision and the other recall. The lexical approach is more conservative, grouping queries into the same cluster only when they share exact term matches, thereby achieving higher precision. In contrast, the semantic approach expands clusters based on term-level semantics, which increases recall but introduces a greater risk of reduced precision within each cluster.
For the sparse representation, we employ the standard TF–IDF vectorization of each query. For the dense representation, we use the [CLS] token’s vector computed with the Sentence-BERT (SBERT) model (reimers-gurevych-2019-sentence). To ensure fair comparison across neural models, we adopt the all-MiniLM-L6-v2 variant of SBERT – a model pre-trained to capture semantic similarities (bowman-etal-2015-large).
@@
(3) $\bar{G}[r(D,\pazocal{C},\pazocal{Q},\theta),\oplus]=\oplus_{i=1}^{K}\operatorname*{G}[r(D,\pazocal{C},\pazocal{Q}_{i},\theta)],$
where $\bar{G}[\cdot]$ represents the Gini coefficient of a set of localised retrievability values, $\oplus\in\{\min,\operatorname*{avg},\max\}$ denotes an aggregation operator. We call this new collection-level measure of document accessibility (Equation 3) by the name Topical-Retrievability, abbreviated as T-Retrievability [^1] The implementation is available at https://github.com/XuejunChang/T-Retrievability .
@@ 3. Experiment Setup
It is worth noting that our proposed localized retrievability scores depend on the granularity of the topical relatedness of queries, i.e., too coarse topics may lead to relevance prior biases, as is expected to be the case with the standard collection-level retrievability (leifRetrievabilityAnevaluationMeasure), whereas too fine-grained topics may fail to capture a collection-level accessibility (Topical_Retrievability2016). Our third research question is thus:
@@
All our experiments corresponding to retrievability analysis and evaluating the relevance of different rankers are conducted based on the MS MARCO dev dataset, which is comprised of 101,093 queries and their associated relevant documents. The underlying collection of documents used is the MS MARCO passage collection, comprised of over 8.8 million passages (bajaj2016ms).
@@
  • BM25 (BM25_RobertsonWJHG94): BM25 is a member of the sparse family of IR models; the term weighting scheme is a combination of the relative importance of terms within documents, term informativeness and document lengths.
  • SPLADE (SPLADE2021): SPLADE, which belongs to the learned-sparse family of rankers, combines the posterior likelihoods of the masked language model (MLM) objective of BERT, and the noise contrastive estimation likelihoods to derive the token weights.
  • TCT-ColBERT (abbreviated in Table 1 as ‘TCT’) (tctcolbert2021): This is a dense end-to-end model distilled from ColBERT that learns effective document encoded vectors. An approximate nearest neighbour search on these embedded representations has been reported to yield better performance than the [CLS] pooling or mean pooling of the ColBERT model. In our experiments, we use the pyterrier_dr library to construct the dense index.
  • BM25¿¿TCT-ColBERT (abbreviated in Table 1 as ‘BM25¿¿TCT’) (tctcolbert2021): the retrieve-and-rerank version of TCT-ColBERT, where we rerank the BM25 top-100 documents with the TCT-ColBERT scores for each query-document pair.
  • BM25¿¿Mono-T5 (abbreviated in Table 1 as ‘Mono-T5’) (monot5): A retrieve-and-rerank class of model, where the top 100 documents retrieved by BM25 are subsequently reranked using a cross-encoder model MonoT5, which assigns relevance scores to each query-document pair.
@@
As evaluation metrics for document exposure bias, we employ the Gini coefficient of the retrievability scores (lower scores indicating fair document exposure). As the measure for relevance, we employ the common measure of nDCG@10 (ndcg10) and MAP (MAP2000).